How much must you deposit in an account that pays 7% semi-annual interest, compounded annually, to have a balance of 4000$ in 15 years?
Answer by addingup(3677)

How much must you deposit in an account that pays 7% semi-annual interest, compounded annually, to have a balance of 4000$ in 15 years?
So you get paid 7% twice per year for a total of 14% per year. Correct? If that's the case:
FV= PV(1+r/n)^nt where
FV is the Future Value
PV is the Present Value
r is the interest rate
n is the number of periods in a year (in your case: 1 period because it's compounded annually)
t= time
4000= PV(1+.14)^15
4000= PV(7.138) Divide both sides by 7.138:
PV= 568.34